OpenStudy (anonymous):

find the absolute value of the complex number 3i-2

OpenStudy (watchmath):

Recall: \(|a+bi|=\sqrt{a^2+b^2}\). Try that first! :)

OpenStudy (anonymous):

how do u solve?

OpenStudy (anonymous):

Using watchmath's formula, the equation 3i-2 would turn into [\sqrt{3^{2}-2^{2}}\]. See if you can solve it from there.

OpenStudy (anonymous):

\[\sqrt{3^{2}-2^{2}}\]

OpenStudy (watchmath):

be careful JetWave600

OpenStudy (watchmath):

a+bi=3i-2 what is a and b here ?

OpenStudy (anonymous):

a=3 b=i ?

OpenStudy (watchmath):

be careful! b is the number before i and a is the number without i

OpenStudy (anonymous):

b=2

OpenStudy (anonymous):

@watchmath

OpenStudy (watchmath):

lets write 3i-2 so it looks like a+bi. We can write it as -2+3i right? now what is a and b?

OpenStudy (anonymous):

3i

OpenStudy (anonymous):

@watchmath

OpenStudy (watchmath):

a+bi=-2+3i. Can you see what is a and b ?

OpenStudy (anonymous):

a=-2 b=3?

OpenStudy (unklerhaukus):

thats right , so what is the modulus (absolute value)?

OpenStudy (anonymous):

how do i find that

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

well ill draw the diagram, maybe then you'll understand where the formula comes from

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):